OpenStudy (anonymous):

PLEASE HELP find the surface area for the given cylinder. Use 3.14/pi: the cylinders base is 6ft and the length is 20ft Answers: a) 180ft b) 720ft c)433ft d)135ft

sammixboo (sammixboo):

The formula to find the SA of a cylinder is A = 2pi*rh + 2pi*r^2 h = height r = radius I am guessing base length = diameter?

sammixboo (sammixboo):

Radius = Half of Diameter

sammixboo (sammixboo):

Kinda confusing. Got a picture?

OpenStudy (anonymous):

yea hold on

sammixboo (sammixboo):

Ok, thanks!

OpenStudy (anonymous):

i was guessing on that answer

sammixboo (sammixboo):

Yes. Our diameter is 6, so our radius would be..?

sammixboo (sammixboo):

Remember. Radius = Half of Diameter

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right?

OpenStudy (wolf1728):

The height is 20 and the radius is 3 right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

@wolf1728

OpenStudy (wolf1728):

The area of the base is 3.14 * 3 * 3 = 28.26 square feet We multiply this by the height which is 20 and we have 565.20 square feet

OpenStudy (anonymous):

is 565.20 the answer?? @wolf1728

OpenStudy (wolf1728):

Let me recheck the math

OpenStudy (anonymous):

because its not coming up on the multiple choice.. the answers are either A) 180ft B) 720ft C)433ft D)135ft

OpenStudy (anonymous):

if you scroll up in the comments i attached a picture of the question with the answer choices

OpenStudy (wolf1728):

lateral cylinder area = 2 * PI * radius * height lateral cylinder area = 376.99 bottom area = 28.26 sq ft total = 405.25 sq ft

OpenStudy (anonymous):

hmm why is the multiple choice not having that answer?? @wolf1728

OpenStudy (anonymous):

oh i got it you have to add the total and bottom area

OpenStudy (wolf1728):

Okay I just thought they wanted the area of the bottom they want the area of the top and bottom so top = 28.26 bottom = 28.26 side = 376.99 TOTAL = 459.25

OpenStudy (wolf1728):

NOPE - the total is 433.51

OpenStudy (wolf1728):

That's the third choice in the multiple choices